# Dynamic Subgraph Matching via Cost-Model-based Vertex Dominance Embeddings (Technical Report)

### v2

@@ 10. Conclusions
In this paper, we formulate and tackle the dynamic subgraph matching (DSM) problem, which continuously monitors the matching subgraph answers over a large-scale dynamic graph. We propose a general framework for efficiently processing DSM queries, based on our carefully-designed vertex dominance embeddings. We also provide an effective degree grouping technique and pruning strategies to facilitate our efficient algorithms of retrieving/maintaining DSM subgraph answers. Most importantly, we devise an effective cost model for guiding the design of vertex embeddings and further enhance the pruning power of our DSM query processing. Through extensive experiments, we evaluate the performance of our DSM approach over real/synthetic dynamic graphs.

### v3

@@ 10. Conclusion
In this paper, we formulate and tackle the continuous subgraph matching (CSM) problem, which continuously monitors the matching subgraph answers over a large-scale dynamic graph. We propose a general framework for efficiently processing CSM queries, based on our carefully-designed DynamIc Vertex DomINance Embedding (DIVINE). We also provide an effective degree grouping technique and pruning strategies to facilitate our efficient algorithms of retrieving/maintaining CSM subgraph answers. Most importantly, we devise an effective cost model for guiding the design of dynamic vertex dominance embeddings and further enhance the pruning power of our CSM query processing. Extensive experiments on real/synthetic graphs confirm our DIVINE query performance.
Acknowledgments
This work was supported by the Natural Science Foundation of China (62272170) and the Shanghai International Joint Lab of Trustworthy Intelligent Software (22510750100). Mingsong Chen is the corresponding author.